OpenStudy (toxicsugar22):

The final buret reading in ml is 31.12 The initial buret reading in ml is 15.72 The volume of titrant in ml is 15.49 what is the moles of NAOH used?

OpenStudy (dan815):

@Somy

OpenStudy (jumperman):

This is actually something I'm in the middle of learning, so I wouldn't know. Sorry.

OpenStudy (anonymous):

Do you have an equation to go with it?

OpenStudy (toxicsugar22):

The final buret reading in ml is 31.12 The initial buret reading in ml is 15.72 The volume of titrant in ml is 15.49

OpenStudy (toxicsugar22):

and they told us to find the moles of NAOH used

OpenStudy (anonymous):

31.12- 15.72 = 15.4 Assuming NaOH is in the burette [seriously? -.-], relative molecular mass of NaOH is 23+16+1 = 40

OpenStudy (anonymous):

volume of titrant is 15.49 ml. They didn't even give you the molarity? Concentration?

OpenStudy (anonymous):

Do you have a molarity of either or the solutions?

OpenStudy (anonymous):

what solutions are you titrating?

OpenStudy (toxicsugar22):

The final buret reading in ml is 31.12 The initial buret reading in ml is 15.72 The volume of titrant in ml is 15.49 what is the moles of NAOH used?

OpenStudy (toxicsugar22):

molarity of standarized NAOH solution is 0.1012 M

OpenStudy (anonymous):

Convert your ml of naoh solution used to litters.

OpenStudy (anonymous):

Sorry i havent done this yet

OpenStudy (anonymous):

then multiply it to the .1012

OpenStudy (anonymous):

\[\frac{moles}{Litters}*Litters=moles\]

OpenStudy (anonymous):

no \[Molarity=\frac{moles}{Litters}\]

OpenStudy (anonymous):

So for this problem \[\frac{.1012 moles}{L}*.001549 L = x moles\]

OpenStudy (anonymous):

check the conversion of ml to litters

OpenStudy (toxicsugar22):

0.0001012 is that the litter

OpenStudy (anonymous):

it looks like you divided the molarity by a thousand when you should have divided the volume reading from the buret

OpenStudy (toxicsugar22):

they also gave me this in the probelm. volume of dilluted vingar used in each titration is 20 mL .02 L

OpenStudy (anonymous):

look at the units you want moles. you have a molarity and a volume of titrant used. if you convert the titrant from ml to L and multiply them together you can get moles by from the information provided by the problem.

OpenStudy (anonymous):

i don't getr it

OpenStudy (triciaal):

I told you this is not fresh in my mind. you first need to balance the equation. you know the mass of one mole. set up the ratio with the volume

OpenStudy (toxicsugar22):

tha is the equation

OpenStudy (toxicsugar22):

Molarity=moles/Litters

OpenStudy (toxicsugar22):

moles/Litters*Litters=moles

OpenStudy (toxicsugar22):

1012moles/L*.001549L=xmoles

OpenStudy (happytalesrebornn):

Sigh... are you kidding me. General Titration? I'm really not willing to help on this one, but @nincompoop is. COUGH COUGH

OpenStudy (aaronq):

what is the moles of NAOH used? Use: \(\sf Molarity_{NaOH} =\dfrac{moles~of~NaOH}{Liters~of~NaOH}\) Molarity=0.1012 M volume=15.49 mL = 0.01549 L plug and solve

OpenStudy (toxicsugar22):

is is .1012/.001549 = 65.33

OpenStudy (triciaal):

no using the formula just given by @aaronq Molarity = moles/volume to get moles you would multiply the molarity by the volume to get moles not divide

OpenStudy (triciaal):



OpenStudy (toxicsugar22):

is it .001567588

OpenStudy (aaronq):

yes it is ! \(\sf Molarity=\dfrac{moles}{L}\rightarrow moles=moles*Molarity=0.01549 L*0.1012 M\) moles=0.001567588 mol

OpenStudy (aaronq):

crap i meant liters*molarity in the formula

OpenStudy (toxicsugar22):

so should i put .001567588 as answer

OpenStudy (toxicsugar22):

asit says how many moles

OpenStudy (aaronq):

nope, you wanna round that up, maybe even use scientific notation

OpenStudy (aaronq):

All your values have 4 sig figs, so you wanna keep 4 sig figs.. \(0.001567588\approx 0.001568\) now put it in scientific notation

OpenStudy (toxicsugar22):

1.567588e-3

OpenStudy (toxicsugar22):

so 1.567*10^-3

OpenStudy (toxicsugar22):

ius it right

OpenStudy (aaronq):

1.568*10^-3

OpenStudy (aaronq):

you have to round up or down depending on the last number you're omitting, in this case that was a 7, so you round to the next number

OpenStudy (toxicsugar22):

ok

OpenStudy (aaronq):

sorry it was a five, 0.001567\(\color{red}5\)88 so it's 1.568..

OpenStudy (toxicsugar22):

thanks

OpenStudy (toxicsugar22):

another question is 15.53 mL = 0.01553 L right

OpenStudy (anonymous):

yes that is correct \[15.53 ml(\frac{1 L}{1000ml})=.01553 L\]